Cary is 8 years older than Dan. In 10 years the sum of their ages will be 102. Find the age of each man now.
Answer by fractalier(6550)

Call their ages NOW, c and d. Then we have
c = d + 8
In 10 years,
(c+10) + (d+10) = 102
Now substitute and get
(d+8+10) + (d+10) = 102
2d + 28 = 102
2d = 74
d = 37
c = 45